OpenStudy (anonymous):

What is the inverse of 3/x+7

OpenStudy (jdoe0001):

\(\bf {\color{red}{ y}}=\cfrac{3}{{\color{blue}{ x}}+7}\qquad inverse\implies {\color{blue}{ x}}=\cfrac{3}{{\color{red}{ y}}+7}\) solve for "y"

OpenStudy (triciaal):

doesn't the inverse mean given what y = then find what x = ? y = 3/(x + 7) 3/y = x + 7 x = 3/y -7